OpenStudy (anonymous):

OKAAAAYY!!!! Now look at this mess how am i suppose to solve this Vincente put his money into a mutual fund, where the amount of money he earned or lost can be found using the equation below, where M is the money Vincente earned or lost and x is time in years. M = x2 - 8x - 109 If Vincente lost $100, how long did he have his money in the mutual fund? 1 year 9 years 8 years 10 years

OpenStudy (anonymous):

8 years

OpenStudy (anonymous):

you mean 9 years

OpenStudy (anonymous):

mainly because you have to Substitute M = -100 into the equation given, and solve for x

OpenStudy (anonymous):

i'll show you my way of doing it 100 = x^2 - 8x - 109 0= x^2 - 8x - 9 0= (x + 1)(x - 9) x + 1 = 0 x - 9 = 0 x= -1 x= 9

OpenStudy (anonymous):

Since x represents time, and time cannot be negative, the solution x = -1 can be discarded

OpenStudy (anonymous):

So Therefore, Vincente left his money in the mutual fund for 9 years

OpenStudy (anonymous):

so are you sure its 8 years

OpenStudy (jojo4eva):

Ok hun look. How are you asking for help and then solving your own question lol. Im so weak right now im about to fall out my chair im laughing so hard lol

OpenStudy (anonymous):

no no im just asking for a response how other people see it @jojo4eva